OpenStudy (anonymous):

The slope of the tangent to the curve y^3x + y^2x^2 = 6 at (2,1)

myininaya (myininaya):

\[(y^3x)'=3y^2y'x+y^3(1)\] remember (y)'=y' \[(y^2x^2)'=2yy'x^2+y^2(2x)\] \[(6)'=0\]

OpenStudy (anonymous):

Then find the slope of the tangent line by taking the first derivative implicitly.

myininaya (myininaya):

all we need is the slope

OpenStudy (anonymous):

myininaya is correct!!!Sorry

myininaya (myininaya):

\[(3y^2y'x+y^3)+(2yy'x^2+y^2(2x))=0\]

OpenStudy (anonymous):

I was giving directions on finding the equation of the tangent line

myininaya (myininaya):

i just plugged in those parts i found above

myininaya (myininaya):

right pross finding the equation of the tangent line is useful too

myininaya (myininaya):

so solve what I have for y' and then plug in x=2 and y=1